There are 64 coins that equal $9.25. How many are dimes and how many are quaters?
Answer by mananth(16949)

let dimes be x
quarters will be 64-x
..
their value = 925 cents
10x+25(64-x)=925
10x+1600-25x=925
-15x=925-1600
-15x=-675
/-15
x=-675/-15
x=45 dimes
..
75 - 45=30 quarters